# Universal Exotic Dynamics in Critical Mesoscopic Systems: Simulating the Square Root of Avogadro’s Number of Spins

### v1

@@ Abstract
We explicitly demonstrate the universality of critical dynamics through unprecedented large-scale GPU-based simulations of two out-of-equilibrium processes, comparing the behavior of spin-$1/2$ Ising and spin-$1$ Blume-Capel models on a square lattice. In the first protocol, a completely disordered system is instantaneously brought into contact with a thermal bath at the critical temperature, allowing it to evolve until the coherence length exceeds $10^{3}$ lattice spacings. Finite-size effects are negligible due to the mesoscopic scale of the lattice sizes studied, with linear dimensions up to $L=2^{22}$ and $2^{19}$ for the Ising and Blume-Capel models, respectively. Our numerical data, and the subsequent analysis, demonstrate a strong dynamic universality between the two models and provide the most precise estimate to date of the dynamic critical exponent for this universality class, $z=2.1676(1)$. In the second protocol, we corroborate the role of the universal ratio of dynamic and static length scales in achieving an exponential acceleration in the approach to equilibrium just above the critical temperature, through a time-dependent variation of the thermal bath temperature. The results presented in this work leverage our CUDA-based numerical code, breaking the world record for the simulation speed of the Ising model.
@@
A particularly fundamental and non-trivial case arises in models whose equilibrium critical properties belong to the same universality class as the Ising ferromagnet. A representative example is the spin-$1$ Blume-Capel model Blume (1966); Capel (1966), specifically in its second-order transition regime Zierenberg et al. (2017). As is well known, the onset of criticality is marked by a divergence of both the correlation length $\xi$ and the correlation time $\tau$. While the former divergence yields singularities in static quantities, the latter manifests notably as critical slowing down. To describe dynamical scaling properties, an additional exponent is required in addition to the static exponents. This so-called dynamic exponent $z$ links the divergence of length and time scales, i.e., $\tau\sim\xi^{z}_{\mathrm{eq}}$ Nightingale and Blöte (1996, 2000); Hasenbusch (2020); Liu et al. (2023) ($\xi_{\mathrm{eq}}$ is the correlation length in equilibrium). In a finite system, $\xi_{\mathrm{eq}}$ is bounded by the linear system size $L$, so that $\tau\sim L^{z}$ at the incipient critical point. The dynamic critical exponent $z$ has been numerically computed to be $z=2.1665(12)$ in two-dimension in the seminal paper by Nightingale and Blöte  Nightingale and Blöte (1996). The more recent theoretical studies based on the nonperturbative renormalization-group Duclut and Delamotte (2017) and the $\varepsilon$-expansion Adzhemyan et al. (2022), suggesting $z\approx 2.15$ and $2.14(2)$, respectively, should also be noted.
@@
where the spins $\sigma_{\mathbf{x}}=\{-1,0,+1\}$ are located in the nodes of an $L\times L$ square lattice with periodic boundary conditions, $\langle\mathbf{x},\mathbf{y}\rangle$ indicates summation over nearest neighbors, and $J>0$ is the ferromagnetic exchange coupling. The parameter $\Delta$ is known as the crystal-field coupling that controls the density of vacancies ($\sigma_{\mathbf{x}}=0$). For $\Delta\rightarrow-\infty$, vacancies are suppressed and the model becomes equivalent to the simple Ising ($\sigma_{\mathbf{x}}=\pm 1$) ferromagnet. The phase boundary of the Blume-Capel model in the crystal field–temperature plane separates the ferromagnetic from the paramagnetic phase Zierenberg et al. (2017). At high temperatures and low crystal fields, the ferromagnetic–paramagnetic transition is a continuous phase transition in the Ising universality class, whereas at low temperatures and high crystal fields, the transition is of first order Blume (1966); Capel (1966). At zero temperature, the ferromagnetic order prevails and the point $(\Delta_{0}=n_{\rm c}J/2,T=0)$, where $n_{\rm c}$ denotes the coordination number, lies on the phase boundary Capel (1966). On the other hand, for zero crystal field, the transition temperature is not exactly known. For the present square-lattice model under study, we provide a high-accuracy estimate $T_{\mathrm{c}}^{\text{(BC)}}(\Delta=0)=1.6935583(5)$; see Appendix B. This result was obtained through a dedicated finite-size scaling analysis, which combines exact results for this universality class Salas and Sokal (2000) with extensive Swendsen-Wang simulations Swendsen and Wang (1987) on systems with linear sizes up to $L\leq 4096$. For comparison, in the case of the simple Ising model (IM), the exact critical temperature used in the simulations below is $T_{\mathrm{c}}^{\text{(IM)}}=2/\log(1+\sqrt{2})$.
@@
Figure 2: Coherence length $\xi$ for a linear size $L=2^{16}$ as a function of time $t$ for both the Ising (main panel) and Blume-Capel (inset) models, following the second thermal protocol.
@@
Figure 3: Estimates of the dynamic critical exponent $z$ for the Ising ($L=2^{22}$) and Blume-Capel ($L=2^{19}$) models obtained via non-local [$\xi$, Eq. (5)] and local [$E$, Eq. (6)] observables, for a varying fitting window ($t_{\text{min}}$: shortest time included in the fit). The main panel displays results from a joint fit for both models (considering data from $\xi$ and $E$ for the Ising model, and only $\xi$ for the Blume-Capel model), as indicated in the legend. The inset shows three sets of data points corresponding to individual fits for each model. We have acceptable values of the fits’ figure of merit $\chi^{2}/\text{dof}$ for all the $t_{\text{min}}$ shown. In both panels, the solid line with the gray-scale background corresponds to $z=2.1665(12)$ Nightingale and Blöte (1996).
@@
$E(t)-E_{\text{eq}}=bt^{-(1/z)}+b^{\prime}t^{-(2/z)}+b^{\prime\prime}t^{-[(\omega+1/\nu)/z]},$ (6)
where again $b$, $b^{\prime}$, and $b^{\prime\prime}$ are non-universal fitting constants. Indeed, for a system of size $L$ in equilibrium, one has $\langle E_{L}\rangle+\sqrt{2}=c_{1}/L+{\cal O}(L^{-2})$, where $c_{1}$ is a known constant Ferdinand and Fisher (1969). Dynamic scaling asserts that we can apply the equilibrium result by replacing $L$ with $\xi(t)$ (noting that amplitudes, such as $c_{1}$, may differ from their equilibrium values). To ensure our determination of $z$ from the energy is independent of long-distance observables, we have further replaced $\xi(t)$ with $t^{1/z}$. Additionally, we have included a correction term, $t^{-[(\omega+1/\nu)/z]}$, to account for scaling corrections in $\xi(t)$, as described in Eq. (5).
Our fitting results are documented in Fig. 3. In particular, the main panel presents the dynamic exponent $z$, which we obtained from a joint fit across three data sets, as indicated in the legend. The estimates for $z$ fall within the range $2.16763(9)-2.1678(5)$ and we report the final result as
@@
Next, we focus on the approach to equilibrium in the paramagnetic phase for $T_{2}\gtrsim T_{\mathrm{c}}$. Specifically, we choose $T_{2}$, such that the equilibrium values of $\xi$ are $\xi_{\text{eq}}(T_{2})=407(1)$ for the Ising model and $405(2)$ for the Blume-Capel model. We then examine the approach to equilibrium of the energy under different thermal protocols, as shown in Fig. 4. The different protocols maintain $T_{1}\approx 0.73T_{\mathrm{c}}$ until $\xi(t,T1)=f\xi_{\text{eq}}(T_{2})$, then the system is instantaneously placed in contact with the thermal bath at $T_{2}$. The direct quench, where $f=0$, is shown in Fig. 4. In the case of direct quench, $E(t)$ approaches equilibrium monotonically from above, whereas the three pre-cooling strategies begin well below the asymptotic equilibrium value. However, $E(t;f=0.54)$ overshoots the equilibrium value of the energy and approaches equilibrium from above, while $E(t;f=0.59)$ and $E(t;f=0.64)$ approach equilibrium from below. An exponential law turns out to aptly fit the late approach to equilibrium
@@
In this work, we have made progress on both fronts, thanks to massive-scale simulations on GPUs. Building on ideas presented in Refs. Romero et al. (2020); Bernaschi et al. (2024), we developed a CUDA code that enabled us to simulate square lattices with side lengths $L=2^{22}$ for the Ising model and $L=2^{19}$ for the Blume-Capel model, reaching coherence lengths well over $10^{3}$ lattice spacings. Thus, while we are taking the limit of large times, the limit of large sizes is certainly taken before. Through this, we have explicitly demonstrated universality by showing that both models are governed by the same dynamic exponent $z$. Along the way, we have obtained the best estimate to date for the dynamic critical exponent, setting a new standard for comparison with field-theoretical computations. Additionally, we considered the pre-cooling strategy from Ref. González-Adalid Pemartín et al. (2021), which achieves an exponential speed-up in the approach to equilibrium at temperatures $T\gtrsim T_{\mathrm{c}}$ by temporarily entering the ferromagnetic phase. Both the Ising and Blume-Capel models exhibit exponential speed-up when the dimensionless ratio of coherence lengths reaches a specific value, a hallmark of a dynamic universality class.
@@ Appendix A Simulation details
The key details of our simulations are presented in Tables 1 and 2. As discussed in the main text and in Appendix B, all simulations related to the Blume-Capel model were conducted at $\Delta=0$.
@@
Table 1: Simulations at a single temperature: All simulations were conducted using our best estimate for $T_{\rm c}$, which is precisely known for the Ising model from the Onsager solution. For the Blume-Capel (BC) model, this estimate was obtained as described in Appendix B. In each case, we report the lattice size $L$, our estimate of $T_{\rm c}$, and the maximum number of Metropolis full-lattice sweeps $t_{\rm max}$. Additionally, we provide the total number of GPU hours, calculated as the sum of the wallclock time multiplied by the number of GPUs used for each run in a set.
@@
Table 2: Simulations at two temperatures: A completely disordered system (i.e., initially at $T=\infty$) began to evolve at a temperature $T_{1}<T_{\rm c}$, until time $t_{\rm switch}$, when the temperature was changed to $T_{\rm 2}$, and the system continued to evolve until the final time step, $t_{\rm max}$. The remaining notational conventions follow those in Table 1. For the Blume-Capel (BC) model, we first conducted an exploratory set of runs on a large timescale ($\sim 2\times 10^{6}$). Based on the resulting statistics, we then performed additional runs, focusing on the cases that required more data. All runs were performed on 4xGH200 GPUs.
@@ Appendix B Critical temperature of the Blume-Capel model
The phase boundary of the model in the crystal field-temperature ($\Delta$, $T$) plane separates the ferromagnetic and paramagnetic phases. The critical line $T_{\mathrm{c}}(\Delta)$ for $0\leq\Delta<\Delta_{\rm t}$, where $\Delta_{\rm t}=1.9660(1)$ Kwak et al. (2015) is the $\Delta$ coordinate of the tricritical point, has been extensively studied in previous numerical works for the square lattice Silva et al. (2006); Malakis et al. (2010); Kwak et al. (2015); Zierenberg et al. (2017). The exact value of the Ising limit for the critical line ($\Delta\to-\infty$), given by $T_{\mathrm{c}}(\Delta\to-\infty)=2/\log(1+\sqrt{2})$, is well known from Onsager’s solution. A consistent trend in the literature indicates that $T_{\mathrm{c}}(\Delta)$ decreases as $\Delta$ increases. This observation motivated our choice of working specifically at $\Delta=0$. By doing so, we aimed to be as far as possible from the Ising $\Delta\rightarrow\infty$ limit while maintaining a high temperature range ***We chose the high-temperature regime because the random number generator used in our CUDA program relies on 32-bit integers Bisson et al. (2025).. In this context, $\Delta=0$ emerged as a reasonable tradeoff, and we set out to improve the accuracy of the known value $T_{c}(\Delta=0)=1.693(3)$ Malakis et al. (2010). To achieve this, we applied finite-size scaling under thermal equilibrium conditions Barber (1983); Cardy (2012); Amit and Martín-Mayor (2005).
We generated equilibrium configurations of the Blume-Capel model at $\Delta=0$ on square lattices with linear sizes $L=2^{n}$ where $5\leq n\leq 12$, using periodic boundary conditions. To achieve equilibration, we combined Swendsen-Wang cluster updates Swendsen and Wang (1987) (restricted to spins with $\sigma_{\mathbf{x}}=\pm 1$) with Metropolis updates that allow for variations in the value of $\sigma_{\mathbf{x}}^{2}$. Our elementary Monte Carlo step (EMCS) in this equilibrium simulation consisted of $10$ full-lattice Metropolis sweeps followed by one iteration of the Swendsen-Wang algorithm. Even for the largest lattice, $L=4096$, we found that $40$ EMCS were sufficient to bring the system to thermal equilibrium. This was verified by comparing simulations starting from a fully random configuration with those starting from the ground state of the Hamiltonian (i.e., $\sigma_{\mathbf{x}}=1$ for all lattice sites ${\mathbf{x}}$). Histogram reweighting was used to extrapolate the results obtained at the simulation’s inverse temperature $\beta$ to neighboring values Falcioni et al. (1982); Ferrenberg and Swendsen (1988). We simulated $100$ independent runs for each lattice size to increase statistical accuracy at the High-Performance Computing cluster CERES of the University of Essex. Each independent run consisted of $4\times 10^{5}$ EMCS (measurements were taken every 40 EMCS). Although it is certainly an overkill, the first 10% of each independent run was discarded to ensure equilibration.
@@
${\cal M}(\mathbf{k})=\sum_{\mathbf{x}}\,{\mathrm{e}}^{\mathrm{i}\mathbf{x}\cdot\mathbf{k}}\sigma_{\mathbf{x}}\,,\quad\chi(\mathbf{k})=\frac{1}{L^{2}}\langle|{\cal M}(\mathbf{k})|^{2}\rangle\,,\quad\mathbf{k}=\frac{2\pi}{L}(n_{x},n_{y})\,,\ n_{x},n_{y}=0,1,\ldots,L-1\,,$ (9)
@@
where we have defined $\chi_{k_{\rm min}}=\chi(2\pi/L,0)=\chi(0,2\pi/L)$. At the critical temperature (for convenience, we use the inverse temperature $\beta=1/T$), $U_{4}$, and $\xi_{2}/L$ reach universal large-$L$ limits, which are known to very high accuracy Salas and Sokal (2000):***Salas and Sokal started from the results of Ref. Di Francesco et al. (1987, 1988) and managed to express both limits in terms of low-dimensional integrals that were evaluated numerically.
$U_{4}^{*}\equiv\lim_{L\to\infty}U_{4}(L,\beta=\beta_{\mathrm{c}})=1.1679229(47)\,,\quad\left[\frac{\xi_{2}}{L}\right]^{*}\equiv\lim_{L\to\infty}\frac{\xi_{2}(L,\beta=\beta_{\text{c}})}{L}=0.9050488292(4)\,.$
The term universal here means that any model in the universality class of the two-dimensional Ising model, and specifically the Blume-Capel model at $\Delta=0$, should exhibit these values for the corresponding quantities on a square lattice. Therefore, we identified the inverse temperatures $\beta_{L,U}$ and $\beta_{L,\xi_{2}/L}$ that satisfy:
@@
Histogram reweighting extrapolations were essential to achieve this goal. As noted in Ref. Amit and Martín-Mayor (2005), dimensionless quantities like $g$ (both $U_{4}$ and $\xi_{2}/L$ are examples) are expected to scale according to $g(L,\beta)=f_{0}[L^{1/\nu}(\beta-\beta_{\mathrm{c}})]+L^{-\omega}f_{1}[L^{1/\nu}(\beta-\beta_{\mathrm{c}})]+\ldots$, where $f_{0}$ and $f_{1}$ are smooth scaling-functions, the dots stand for sub-dominant corrections to scaling, and $\nu=1$ for this universality class. Defining $\beta_{L,g}$ through $g(L,\beta_{L,g})=f_{0}(0)$, we obtain the scaling form
@@
where $\mathcal{A}$ is a scaling amplitude and the dots indicate sub-leading corrections to scaling. Interestingly, the corrections-to-scaling exponent $\omega$ is not well known for the two-dimensional Ising model universality class. In practice, one often uses $\omega=7/4$, which corresponds to the leading analytic corrections to scaling, dominating over the irrelevant renormalization-group eigenvalues. In fact, both $\beta_{L,U}$ and $\beta_{L,\xi_{2}/L}$, as shown in Fig. 5, approach their large-$L$ limits faster than predicted by the scaling form in Eq. (12). For all our results with $L\geq 64$, we observed compatibility with a constant value. As a result, we fit the data to a constant, as the statistical errors for $\beta_{L,U}$ and $\beta_{L,\xi_{2}/L}$ decrease significantly with increasing $L$. Thus, the fit essentially yields the value for $L=4096$. This process led to the estimate of $T_{\mathrm{c}}(\Delta=0)$, as quoted in the main text.

### v2

@@ Abstract
We explicitly demonstrate the universality of critical dynamics through unprecedented large-scale GPU-based simulations of two out-of-equilibrium processes, comparing the behavior of spin-1/21/21 / 2 Ising and spin-111 Blume-Capel models on a square lattice. In the first protocol, a completely disordered system is instantaneously brought into contact with a thermal bath at the critical temperature, allowing it to evolve until the coherence length exceeds $10^{3}$ lattice spacings. Finite-size effects are negligible due to the mesoscopic scale of the lattice sizes studied, with linear dimensions up to $L=2^{22}$ and $2^{19}$ for the Ising and Blume-Capel models, respectively. Our numerical data, and the subsequent analysis, demonstrate a strong dynamic universality between the two models and provide the most precise estimate to date of the dynamic critical exponent for this universality class, $z=2.1676(1)$. In the second protocol, we corroborate the role of the universal ratio of dynamic and static length scales in achieving an exponential acceleration in the approach to equilibrium just above the critical temperature, through a time-dependent variation of the thermal bath temperature. The results presented in this work leverage our CUDA-based numerical code, breaking the world record for the simulation speed of the Ising model.
@@
A particularly fundamental and non-trivial case arises in models whose equilibrium critical properties belong to the same universality class as the Ising ferromagnet. A representative example is the spin-111 Blume-Capel model Blume (1966); Capel (1966), specifically in its second-order transition regime Zierenberg et al. (2017). As is well known, the onset of criticality is marked by a divergence of both the correlation length $\xi$ and the correlation time $\tau$. While the former divergence yields singularities in static quantities, the latter manifests notably as critical slowing down. To describe dynamical scaling properties, an additional exponent is required in addition to the static exponents. This so-called dynamic exponent $z$ links the divergence of length and time scales, i.e., $\tau\sim\xi^{z}_{\mathrm{eq}}$ Nightingale and Blöte (1996, 2000); Hasenbusch (2020); Liu et al. (2023) ($\xi_{\mathrm{eq}}$ is the correlation length in equilibrium). In a finite system, $\xi_{\mathrm{eq}}$ is bounded by the linear system size $L$, so that $\tau\sim L^{z}$ at the incipient critical point. The dynamic critical exponent $z$ has been numerically determined to be $z=2.1665(12)$ in two dimensions in the seminal work by Nightingale and Blöte Nightingale and Blöte (1996), and was later shown to be universal with respect to the underlying lattice structure Wang and Hu (1997). The more recent theoretical studies based on the nonperturbative renormalization-group Duclut and Delamotte (2017) and the $\varepsilon$-expansion Adzhemyan et al. (2022), suggesting $z\approx 2.15$ and 2.14(2)2.14(2)2.14 ( 2 ), respectively, should also be noted.
@@
where the spins $\sigma_{\mathbf{x}}=\{-1,0,+1\}$ are located in the nodes of an $L\times L$ square lattice with periodic boundary conditions, $\langle\mathbf{x},\mathbf{y}\rangle$ indicates summation over nearest neighbors, and $J>0$ is the ferromagnetic exchange coupling. The parameter $\Delta$ is known as the crystal-field coupling that controls the density of vacancies ($\sigma_{\mathbf{x}}=0$). For $\Delta\rightarrow-\infty$, vacancies are suppressed and the model becomes equivalent to the simple Ising ($\sigma_{\mathbf{x}}=\pm 1$) ferromagnet. The phase boundary of the Blume-Capel model in the crystal field–temperature plane separates the ferromagnetic from the paramagnetic phase Zierenberg et al. (2017). At high temperatures and low crystal fields, the ferromagnetic–paramagnetic transition is a continuous phase transition in the Ising universality class, whereas at low temperatures and high crystal fields, the transition is of first order Blume (1966); Capel (1966). At zero temperature, the ferromagnetic order prevails and the point $(\Delta_{0}=n_{\rm c}J/2,T=0)$, where $n_{\rm c}$ denotes the coordination number, lies on the phase boundary Capel (1966). On the other hand, for zero crystal field, the transition temperature is not exactly known. For the present square-lattice model under study, we provide a high-accuracy estimate $T_{\mathrm{c}}^{\text{(BC)}}(\Delta=0)=1.6935583(5)$; see Appendix B. This result was obtained through a dedicated finite-size scaling analysis, which combines exact results for this universality class Di Francesco et al. (1987, 1988); Salas and Sokal (2000); Caselle et al. (2002) with extensive Swendsen-Wang simulations Swendsen and Wang (1987) on systems with linear sizes up to $L\leq 4096$. For comparison, in the case of the simple Ising model (IM), the exact critical temperature used in the simulations below is $T_{\mathrm{c}}^{\text{(IM)}}=2/\log(1+\sqrt{2})$.
@@
Figure 2: Coherence length $\xi$ for a linear size $L=2^{16}$ as a function of time $t$ for both the Ising (main panel) and Blume-Capel (inset) models, following the second thermal protocol. $\xi(t)$ appears to be almost $f$-independent in this case.
@@
Figure 3: Estimates of the dynamic critical exponent $z$ for the Ising ($L=2^{22}$) and Blume-Capel ($L=2^{19}$) models obtained via non-local [$\xi$, Eq. (15)] and local [$E$, Eq. (16)] observables, for a varying fitting window ($t_{\text{min}}$: shortest time included in the fit). The main panel displays results from a joint fit for both models (considering data from $\xi$ and $E$ for the Ising model, and only $\xi$ for the Blume-Capel model), as indicated in the legend. The inset shows three sets of data points corresponding to individual fits for each model. We have acceptable values of the fits’ figure of merit $\chi^{2}/\text{dof}$, where ${\rm dof}$ stands for the number of degrees of freedom, for all the $t_{\text{min}}$ shown. In both panels, the solid line with the gray-scale background corresponds to $z=2.1665(12)$ Nightingale and Blöte (1996).
@@
$E(t)-E_{\text{eq}}=bt^{-(1/z)}+b^{\prime}t^{-(2/z)}+b^{\prime\prime}t^{-[(\omega+1)/z]},$ (6)
where again $b$, $b^{\prime}$, and $b^{\prime\prime}$ are non-universal fitting constants. Indeed, for a system of size $L$ in equilibrium, one has $\langle E_{L}\rangle+\sqrt{2}=c_{1}/L+{\cal O}(L^{-2})$, where $c_{1}$ is a known constant Ferdinand and Fisher (1969). Dynamic scaling asserts that we can apply the equilibrium result by replacing $L$ with $\xi(t)$ (noting that amplitudes, such as $c_{1}$, may differ from their equilibrium values). To ensure our determination of $z$ from the energy is independent of long-distance observables, we have further replaced $\xi(t)$ with $t^{1/z}$. Additionally, we have included a correction term, $t^{-[(\omega+1)/z]}$, to account for scaling corrections in $\xi(t)$, as described in Eq. (15) that we have inverted as $\xi=B_{0}\,t^{1/z}(1+B_{1}/t^{\omega/z}+\cdots)$ where $B_{0}$ and $B_{1}$ are scaling amplitudes.
Our fitting results are documented in Fig. 3. In particular, the main panel presents the dynamic exponent $z$, which we obtained from a joint fit across three data sets, as indicated in the legend. The estimates for $z$ fall within the range 2.16763(9)2.1678(5)2.16763(9)-2.1678(5)2.16763 ( 9 ) - 2.1678 ( 5 ) and we report the final result as
@@
Next, we focus on the approach to equilibrium in the paramagnetic phase for $T_{2}\gtrsim T_{\mathrm{c}}$. Specifically, we choose $T_{2}$, such that the equilibrium values of $\xi$ are $\xi_{\text{eq}}(T_{2})=407(1)$ for the Ising model and 405(2)405(2)405 ( 2 ) for the Blume-Capel model. We then examine the approach to equilibrium of the energy under different thermal protocols, as shown in Fig. 4. The different protocols maintain $T_{1}\approx 0.73T_{\mathrm{c}}$ until $\xi(t,T1)=f\xi_{\text{eq}}(T_{2})$, then the system is instantaneously placed in contact with the thermal bath at $T_{2}$. The direct quench, where $f=0$, is shown in Fig. 4. In the case of direct quench, $E(t)$ approaches equilibrium monotonically from above, whereas the three pre-cooling strategies begin well below the asymptotic equilibrium value. However, $E(t;f=0.54)$ overshoots the equilibrium value of the energy and approaches equilibrium from above, while $E(t;f=0.59)$ and $E(t;f=0.64)$ approach equilibrium from below. An exponential law turns out to aptly fit the late approach to equilibrium
@@
In this work, we have made progress on both fronts, thanks to massive-scale simulations on GPUs. Building on ideas presented in Refs. Romero et al. (2020); Bernaschi et al. (2024), we developed a CUDA code that enabled us to simulate square lattices with side lengths $L=2^{22}$ for the Ising model and $L=2^{19}$ for the Blume-Capel model, reaching coherence lengths well over $10^{3}$ lattice spacings. Thus, while we are taking the limit of large times, the limit of large sizes is certainly taken before. Through this, we have explicitly demonstrated universality by showing that both models are governed by the same dynamic exponent $z$. Along the way, we have obtained the best estimate to date for the dynamic critical exponent, setting a new standard for comparison with field-theoretical computations. Additionally, we considered the pre-cooling strategy from Ref. González-Adalid Pemartín et al. (2021), which achieves an exponential speed-up in the approach to equilibrium at temperatures $T\gtrsim T_{\mathrm{c}}$ by temporarily entering the ferromagnetic phase. Both the Ising and Blume-Capel models exhibit exponential speed-up when the dimensionless ratio of coherence lengths reaches a specific value, a hallmark of a dynamic universality class. We anticipate that the numerical approach developed in this study may be adapted to help resolve ongoing debates about the universality classes of other nonequilibrium systems, including those encountered in active matter Chen et al. (2025).
@@ Appendix A Simulation details
The key details of our simulations are presented in Tables 1 and 2. As discussed in the main text and in Sec. B, all simulations related to the Blume-Capel model were conducted at $\Delta=0$.
@@
Table 1: Simulations at a single temperature: All simulations were conducted using our best estimate for $T_{\rm c}$, which is precisely known for the Ising model from the Onsager solution. For the Blume-Capel model, this estimate was obtained as described in Sec. B. In each case, we report the lattice size $L$, our estimate of $T_{\rm c}$, and the maximum number of Metropolis full-lattice sweeps $t_{\rm max}$. Additionally, we provide the total number of GPU hours, calculated as the sum of the wallclock time multiplied by the number of GPUs used for each run in a set.
@@
Table 2: Simulations at two temperatures: A completely disordered system (i.e., initially at $T=\infty$) began to evolve at a temperature $T_{1}<T_{\rm c}$, until time $t_{\rm switch}$, when the temperature was changed to $T_{\rm 2}$, and the system continued to evolve until the final time step, $t_{\rm max}$. The remaining notational conventions follow those in Table 1. For the Blume-Capel model, we first conducted an exploratory set of runs on a large timescale ($\sim 2\times 10^{6}$). Based on the resulting statistics, we then performed additional runs, focusing on the cases that required more data. All runs were performed on 4xGH200 GPUs.
@@ Appendix B Critical temperature of the Blume-Capel model
The phase boundary of the model in the crystal field-temperature ($\Delta$, $T$) plane separates the ferromagnetic and paramagnetic phases. The critical line $T_{\mathrm{c}}(\Delta)$ for $0\leq\Delta<\Delta_{\rm t}$, where $\Delta_{\rm t}=1.9660(1)$ Kwak et al. (2015) is the $\Delta$ coordinate of the tricritical point, has been extensively studied in previous numerical works for the square lattice Silva et al. (2006); Malakis et al. (2010); Kwak et al. (2015); Zierenberg et al. (2017). The exact value of the Ising limit for the critical line ($\Delta\to-\infty$), given by $T_{\mathrm{c}}(\Delta\to-\infty)=2/\log(1+\sqrt{2})$, is well known from Onsager’s solution. A consistent trend in the literature indicates that $T_{\mathrm{c}}(\Delta)$ decreases as $\Delta$ increases. This observation motivated our choice of working specifically at $\Delta=0$. By doing so, we aimed to be as far as possible from the Ising $\Delta\rightarrow\infty$ limit while maintaining a high temperature range. We chose the high-temperature regime because the random number generator used in our CUDA program relies on 32-bit integers Bisson et al. (2025). In this context, $\Delta=0$ emerged as a reasonable trade-off, and we set out to improve the accuracy of the known value $T_{c}(\Delta=0)=1.693(3)$ Malakis et al. (2010). To achieve this, we applied finite-size scaling under thermal equilibrium conditions Barber (1983); Cardy (2012); Amit and Martín-Mayor (2005).
We generated equilibrium configurations of the Blume-Capel model at $\Delta=0$ on square lattices with linear sizes $L=2^{n}$ where $5\leq n\leq 12$, using periodic boundary conditions. To achieve equilibration, we combined Swendsen-Wang cluster updates Swendsen and Wang (1987) (restricted to spins with $\sigma_{\mathbf{x}}=\pm 1$) with Metropolis updates that allow for variations in the value of $\sigma_{\mathbf{x}}^{2}$. Our elementary Monte Carlo step (EMCS) in this equilibrium simulation consisted of 101010 full-lattice Metropolis sweeps followed by one iteration of the Swendsen-Wang algorithm. Even for the largest lattice, $L=4096$, we found that 404040 EMCS were sufficient to bring the system to thermal equilibrium. This was verified by comparing simulations starting from a fully random configuration with those starting from the ground state of the Hamiltonian (i.e., $\sigma_{\mathbf{x}}=1$ for all lattice sites ${\mathbf{x}}$). Histogram reweighting was used to extrapolate the results obtained at the simulation’s inverse temperature $\beta$ to neighboring values Falcioni et al. (1982); Ferrenberg and Swendsen (1988). We simulated 100100100 independent runs for each lattice size to increase statistical accuracy at the High-Performance Computing cluster CERES of the University of Essex. Each independent run consisted of $4\times 10^{5}$ EMCS (measurements were taken every 40 EMCS). Although it is certainly an overkill, the first 10% of each independent run was discarded to ensure equilibration.
@@
${\cal M}(\mathbf{k})=\sum_{\mathbf{x}}\,{\mathrm{e}}^{\mathrm{i}\mathbf{x}\cdot\mathbf{k}}\sigma_{\mathbf{x}}\,,\quad\chi(\mathbf{k})=\frac{1}{L^{2}}\langle|{\cal M}(\mathbf{k})|^{2}\rangle\,,\quad$
$\mathbf{k}=\frac{2\pi}{L}(n_{x},n_{y})\,,\ n_{x},n_{y}=0,1,\ldots,L-1\,,$ (9)
@@
[…]
$U_{4}^{*}\equiv\lim_{L\to\infty}U_{4}(L,\beta=\beta_{\mathrm{c}})=1.1679229(47)\,,\quad$
$\left[\frac{\xi_{2}}{L}\right]^{*}\equiv\lim_{L\to\infty}\frac{\xi_{2}(L,\beta=\beta_{\text{c}})}{L}=0.9050488292(4)\,.$
We note here that Salas and Sokal started from the results of Refs. Di Francesco et al. (1987, 1988) and managed to express both limits in terms of low-dimensional integrals that were evaluated numerically. The term universal here means that any model in the universality class of the two-dimensional Ising model, and specifically the Blume-Capel model at $\Delta=0$, should exhibit these values for the corresponding quantities on a square lattice. Therefore, we identified the inverse temperatures $\beta_{L,U}$ and $\beta_{L,\xi_{2}/L}$ that satisfy:
@@
Histogram reweighting extrapolations were essential to achieve this goal. As noted in Ref. Amit and Martín-Mayor (2005), dimensionless quantities like $g$ (both $U_{4}$ and $\xi_{2}/L$ are examples) are expected to scale according to $g(L,\beta)=f_{0}[L^{1/\nu}(\beta-\beta_{\mathrm{c}})]+L^{-\omega}f_{1}[L^{1/\nu}(\beta-\beta_{\mathrm{c}})]+\cdots$, where $f_{0}$ and $f_{1}$ are smooth scaling-functions, the dots stand for sub-dominant corrections to scaling, and $\nu=1$ for this universality class. Defining $\beta_{L,g}$ through $g(L,\beta_{L,g})=f_{0}(0)$, we obtain the scaling form
@@
where $\mathcal{A}$ is a scaling amplitude and the dots indicate sub-leading corrections to scaling. The corrections to scaling in the two-dimensional Ising model universality class have been elucidated using tools from conformal field theory Caselle et al. (2002).
[…]